# Multilevel neural simulation-based inference

### v1

@@ 4 Theory
We now present our main theoretical results. Theorem 1 expresses the variance of each of the terms in the telescoping sum approximation (see 2) as a function of the number of simulations per level, and the magnitude of the difference in generators between consecutive levels. We say that $\mu$ is log-concave if it has a density of the form $\exp(-\psi(z))$ for some convex function $\psi:\mathcal{Z}\rightarrow\mathbb{R}$. We recall that for $r\in[1,\infty]$, $d,d^{\prime}\in\mathbb{N}$ and a non-empty, open, connected set $\mathcal{Z}\subseteq\mathbb{R}^{d}$, the space of vector-valued $r$-integrable functions with respect to a probability distribution $\mu$ is given by $L^{r}(\mu):=\{g:\mathcal{Z}\rightarrow\mathbb{R}^{d^{\prime}}:\|g\|_{L^{r}(\mu)}:=(\int_{\mathcal{Z}}\|g(x)\|_{2}^{r}\mu(dx))^{\nicefrac{{1}}{{r}}}<\infty\}$. For $\tau\in\mathbb{N}$, the corresponding Sobolev space of vector-valued functions of smoothness $\tau$ is given by $W^{\tau,r}(\mu):=\{g:\mathcal{Z}\rightarrow\mathbb{R}^{d^{\prime}}:\|g\|_{W^{\tau,r}(\mu)}=(\sum_{|\alpha|\leq\tau}\|D^{\alpha}g\|_{L^{r}(\mu)}^{r})^{\nicefrac{{1}}{{r}}}<\infty\}$, where for a multi-index $\alpha\in\mathbb{N}^{d}$, $D^{\alpha}$ is the weak derivative operator corresponding to $\alpha$. Finally, we recall that a function $g$ is locally $K_{\text{Lip}}$-smooth if its gradient is locally Lipschitz continuous with Lipschitz constant $K_{\text{Lip}}>0$; i.e. for all $z,z^{\prime}$ in some open set of $\mathcal{Z}$, we have that $\|\nabla g(z)-\nabla g(z^{\prime})\|_{2}\leq K_{\text{Lip}}\|z-z^{\prime}\|_{2}$.
@@
    • NLE: $x,\theta\mapsto\log q^{\text{\text{NLE}}}_{\phi}(x|\theta)$ is locally $K_{\text{Lip}}(\phi)-$smooth and satisfies the growth condition $\|\nabla_{x}\log q^{\text{\text{NLE}}}_{\phi}(x|\theta)\|_{2}\leq K_{\text{grow}}(\phi)(\|(x,\theta)\|_{2}+1)$ for some $K_{\text{Lip}}(\phi),K_{\text{grow}}(\phi)>0$.
    • NPE: $x_{1},\ldots,x_{m},\theta\mapsto\log q^{\text{\text{NPE}}}_{\phi}(\theta|x_{1},\ldots,x_{m})$ is locally $K_{\text{Lip}}(\phi)-$smooth and satisfies the growth condition $\|\nabla_{\theta}\log q^{\text{\text{NLE}}}_{\phi}(\theta|x_{1},\ldots,x_{m})\|_{2}\leq K_{\text{grow}}(\phi)(\|(x_{1},\ldots,x_{m},\theta)\|_{2}+1)$ for some $K_{\text{Lip}}(\phi),K_{\text{grow}}(\phi)>0$.
@@
$\text{and}\;\text{Var}\big{[}h_{l}(\phi)\big{]}$ $\leq\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U}^{m})}+1\right)$
See Section B.1 for the proof, and we emphasise that the variance is over both parameters $\theta$ and noise $u_{1:m}$. (A1) requires log-concavity, which is a strong condition, but this is only required of the prior and the base measure. However, in SBI these tend to be simple distributions such as Gaussians or uniforms, which satisfy this assumption; see Saumard and Wellner (2014) for how to verify log-concavity in practice. (A2) is relatively mild: it asks that $G^{0},\ldots,G^{L}$ have at least one derivative in $\theta$ and $u$, and for these generators and their derivatives to have a fourth moment. This will hold when the simulators are used to define sufficiently well-behaved data-generating processes, but will be violated for sufficiently heavy-tailed distributions (e.g. the Cauchy). Finally, (A3) is mild; it holds when the gradient of the log conditional density estimator is Lipschitz continuous, which is for example the case when the conditional density is twice continuously differentiable and the Hessian is bounded (see e.g. Lemma 2.3 of Wright and Recht (2022)). It also holds for models such as mixtures of Gaussians, which are used in mixture density networks and for normalising flows with sufficiently regular transformations. The key implication of this result is a bound on the variance of our objective:
$\text{Var}\left[{\ell}_{\text{MLMC}}(\phi)\right]$ $\leq\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}^{4}+1\right)$
$\qquad+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U}^{m})}+1\right)$ (3)
In order to make this bound small, we need to make each term small. Typically, $\|G^{0}\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}$ will be large, but this will be counter-balanced by taking $n_{0}$ to be large. For the higher-fidelity levels, the number of samples $n_{l}$ will typically be smaller, but this will be counter-balanced by the fact that $\|G^{l}-G^{l-1}\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}$ is small whenever $G^{l-1}$ is a good approximation of $G^{l}$. As we now show, we can directly use this bound to get an indication of how to select $n_{0},\ldots,n_{L}$.
@@
Suppose the assumptions of Theorem 1 hold. Then, the values of $n_{0},\ldots,n_{L}$ which minimise the upper bound on $\text{Var}[{\ell}_{\text{MLMC}}(\phi)]$ in 3 given a fixed computational budget; i.e. for $\text{Cost}({\ell}_{\text{MLMC}}(\phi);n_{0},\ldots,n_{L})\leq C_{\text{budget}}$ for some $C_{\text{budget}}>0$ are given by
$n^{\star}_{0}\propto\frac{C_{\text{budget}}}{\sqrt{C_{0}}}\sqrt{\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}^{4}+1},\quad n^{\star}_{l}\propto\frac{C_{\text{budget}}}{\sqrt{C_{l}+C_{l+1}}}\sqrt{\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U}^{m})}+1}.$
See Section B.2 for the proof. The denominator in these expressions demonstrates that the more expensive the simulator is at level $l$, the smaller we should make the number of simulations at that level. We note that this result provides useful intuition on how to select the number of sample per level, but may be hard to use in practice since it requires quantities which are often unknown. For example, computing Sobolev norms can be challenging, and the results implicitly depends on $\phi$ through the constants $K_{0}(\phi),\ldots,K_{L}(\phi)$ of Theorem 1 (which are unlikely to be tight). This is a common limitation of MLMC theory; see Section 2 and 3 of Giles (2015) for a discussion.
@@ 5 Numerical Experiments
We compare the performance of our multilevel version of NLE and NPE with the MLMC loss, termed ML-NLE and ML-NPE, respectively, against their standard counterpart with the MC loss. We use the sbi library (Tejero-Cantero et al., 2020) implementation for NLE and NPE , see Appendix C for the details. The code to reproduce our experiments is available at https://github.com/yugahikida/multilevel-sbi.
@@ 5.1 The g-and-k distribution: an illustrative example
We first consider the g-and-k distribution (Prangle, 2020) as an illustrative example. This is a very flexible univariate distribution that is defined via its quantile function and has four parameters, controlling the mean, variance, skewness, and kurtosis respectively, making it challenging for SBI methods. It does not typically have a low-fidelity simulator, so we construct one through a Taylor approximation of the quantile function, see Section C.1 for details. This makes it a slightly contrived example, but the fact that the g-and-k allows for an efficient approximation of the likelihood will make it particularly convenient to study the performance of NLE-based methods.
@@
We fix the number of low-fidelity samples ($n_{0}=10^{4}$ for ML-NLE and $n_{0}=10^{3}$ for ML-NPE) and vary the number of high-fidelity samples $n_{1}$ to asses the improvement in performance of our methods as $n_{1}$ increases. For ML-NLE, we compute the Kullback-Leibler divergence (KLD) between the estimated conditional density and a numerical approximation of the likelihood. For ML-NPE, we use the negative log-posterior density (NLPD) of the true $\theta$ under the estimated posterior density as the metric. The results in Figure 2(a) and 2(b) show that the multilevel versions of NLE and NPE perform better than their standard counterparts (with MC loss) using just a fraction of the high-fidelity samples. Unsurprisingly, the performance of MLE-NLE and ML-NPE improves as $n_{1}$ increases.
In Figure 2(c), we show an example of the NLE densities learned, with additional examples in Section C.1. Our ML-NLE method is able to approximate the almost exact g-and-k density the best. The coverage plot in Figure 2(d) shows that ML-NPE yields slightly conservative posteriors as opposed to the overconfident posteriors obtained from NPE trained on either all low- or all high-fidelity data.
@@ 5.2 Ornstein–Uhlenbeck process: a popular financial model
Our next experiment involves the Ornstein-Uhlenbeck (OU) process—a stochastic differential equation model commonly used in financial analysis (Minenna, 2003), which in our case outputs a $100$-dimensional Markovian time-series and has three parameters. The process is known to converge to a stationary Gaussian distribution, which we use as the low-fidelity simulator, see Section C.3. The example is reproduced from Krouglova et al. (2025), who used it for benchmarking their transfer learning approach to NPE (TL-NPE). TL-NPE first trains an NPE network on low-fidelity simulations, and then uses a small set of high-fidelity data to refine the network parameters until a stopping criterion is met. We implement TL-NPE with $n_{0}=11,000$ and $n_{1}=1000$, using 10% of data as validation set for their stopping criterion, and compare against ML-NPE with $n_{0}=10,000$ and $n_{1}=1000$. These values are picked so as to keep the simulation budget same for both the methods.
We compare ML-NPE with TL-NPE for different choices of its hyperparameter in Figure 3. TL-NPE’s hyperparameter is the number of epochs until the stopping criterion is activated, termed patience, which we let take values in $\{1,20,10^{2},10^{3}\}$. We use two metrics for evaluation: the NLPD of the true $\theta$ under the estimated posterior, and the KLD between the posterior obtained using either TL-NPE or ML-NPE and a reference NPE posterior trained using $n=100,000$ samples. We observe that TL-NPE’s performance is quite sensitive to the choice of the patience value, and the authors of that work do not provide guidance on how to select this in practice (their code fixes this to $20$ with no clear justification). When the number of high-fidelity samples are low ($n_{1}=10$), its performance degrades as patience increases, as a higher patience means TL-NPE is overfitting to the few high-fidelity samples, losing the information gained from training on the low-fidelity data. In contrast, ML-NPE achieves better performance especially when $n_{1}=10$ and does not have any hyperparameter to tune, making it easier to use in practice.
@@
Figure 3: NLPD and KLD for ML-NPE (ours) and TL-NPE (Krouglova et al., 2025) using different patience values (hyperparameter of TL-NPE) over $20$ runs. ML-NPE performs better in the limited high-fidelity data scenario ($n_{1}=10$) (a), whilst being competitive when $n_{1}=100$ (b). TL-NPE shows sensitivity towards its hyperparameter, while ML-NPE does not require any hyperparameter.
@@ 5.3 Toggle-switch model: a Systems Biology example
Figure 4: Boxplots of MMD ($\downarrow$) across 5000 true parameter values for NLE and ML-NLE.
We now consider the toggle-switch model (Bonassi et al., 2011; Bonassi and West, 2015). This model describes the interaction between two genes over time, and has seven parameters and a scalar observation at the end of a time interval. Simulators typically use time-discretisation, and the total number of time-steps $T$ acts as a fidelity parameter: running the model with large $T$ incurs larger computational cost but leads to accurate simulations, while smaller $T$ leads to cheap but inaccurate samples. Thus, this model illustrates a setting with more than two fidelity levels (by taking $T$ to be more than two values)—a situation straightforward to handle with our multilevel approach but not with TL-NPE (Krouglova et al., 2025).
We take $T_{0}=50$, $T_{1}=80$, and $T_{2}=300$ to be the number of steps for the three fidelity levels, with $n_{0}=10^{4}$, $n_{1}=500$, and $n_{2}=300$. Note that in this case each fidelity level has a different base measure of dimension $12T+$ 1; however, there is still sufficient seed-matching (see Section C.4 for details) and hence variance reduction thanks to MLMC. We compare our ML-NLE method with the standard NLE trained on samples from either the low- ($n=12,060$), the medium- ($n=7537$), or the high-fidelity simulator ($n=2010$). The number of training data $n$ in each case is selected so as to match the total computational cost of simulation between ML-NLE and NLE, see Section C.4. Figure 4 reports the maximum mean discrepancy (MMD) (Gretton et al., 2012) between $500$ samples from the learned conditional densities and $500$ samples from the high-fidelity simulator across. We observe that ML-NLE performs better than all the NLE baselines for the same computational cost.

### v2

@@ 4 Theory
We now present our main theoretical results. Theorem˜1 expresses the variance of each of the terms in the telescoping sum approximation (see ˜2) as a function of the number of simulations per level, and the magnitude of the difference in generators between consecutive levels.
We say that $\mu$ is log-concave if it has a density of the form $\exp(-\psi(z))$ for some convex function $\psi:\mathcal{Z}\rightarrow\mathbb{R}$. We recall that for $r\in[1,\infty]$, $d,d^{\prime}\in\mathbb{N}$ and a non-empty, open, connected set $\mathcal{Z}\subseteq\mathbb{R}^{d}$, the space of vector-valued $r$-integrable functions with respect to a probability distribution $\mu$ is given by $L^{r}(\mu):=\{g:\mathcal{Z}\rightarrow\mathbb{R}^{d^{\prime}}:\|g\|_{L^{r}(\mu)}:=(\int_{\mathcal{Z}}\|g(x)\|_{2}^{r}\mu(dx))^{\nicefrac{{1}}{{r}}}<\infty\}$. For $\tau\in\mathbb{N}$, the corresponding Sobolev space of vector-valued functions of smoothness $\tau$ is given by $W^{\tau,r}(\mu):=\{g:\mathcal{Z}\rightarrow\mathbb{R}^{d^{\prime}}:\|g\|_{W^{\tau,r}(\mu)}=(\sum_{|\alpha|\leq\tau}\|D^{\alpha}g\|_{L^{r}(\mu)}^{r})^{\nicefrac{{1}}{{r}}}<\infty\}$, where for a multi-index $\alpha\in\mathbb{N}^{d}$, $D^{\alpha}$ is the weak derivative operator corresponding to $\alpha$. Finally, we recall that a function $g$ is locally $K_{\text{Lip}}$-smooth if its gradient is locally Lipschitz continuous with Lipschitz constant $K_{\text{Lip}}>0$; i.e. for all $z,z^{\prime}$ in some open set of $\mathcal{Z}$, we have that $\|\nabla g(z)-\nabla g(z^{\prime})\|_{2}\leq K_{\text{Lip}}\|z-z^{\prime}\|_{2}$.
@@
    • NLE: $x,\theta\mapsto\log q^{\text{\text{NLE}}}_{\phi}(x|\theta)$ is continuously differentiable, locally $K_{\text{Lip}}(\phi)$-smooth and satisfies the growth condition $\|\nabla_{x}\log q^{\text{\text{NLE}}}_{\phi}(x|\theta)\|_{2}\leq K_{\text{grow}}(\phi)(\|(x,\theta)\|_{2}+1)$ for some $K_{\text{Lip}}(\phi),K_{\text{grow}}(\phi)>0$.
    • NPE: $x_{1},\ldots,x_{m},\theta\mapsto\log q^{\text{\text{NPE}}}_{\phi}(\theta|x_{1},\ldots,x_{m})$ is continously differentiable, locally $K_{\text{Lip}}(\phi)$-smooth and satisfies the growth condition $\|\nabla_{\theta}\log q^{\text{\text{NLE}}}_{\phi}(\theta|x_{1},\ldots,x_{m})\|_{2}\leq K_{\text{grow}}(\phi)(\|(x_{1},\ldots,x_{m},\theta)\|_{2}+1)$ for some $K_{\text{Lip}}(\phi),K_{\text{grow}}(\phi)>0$.
@@
$\text{and}\;\text{Var}\big{[}h_{l}(\phi)\big{]}$ $\leq\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U}^{m})}+1\right).$
See Section˜A.1 for the proof, and we emphasise that the variance is over both parameters $\theta$ and noise $u_{1:m}$. (A1) requires log-concavity, which is a strong condition, but this is only required of the prior and the base measure. However, in SBI these tend to be simple distributions such as Gaussians or uniforms, which satisfy this assumption; see Saumard and Wellner (2014) for how to verify log-concavity in practice. (A2) is relatively mild: it asks that $G^{0},\ldots,G^{L}$ have at least one derivative in $\theta$ and $u$, and for these generators and their derivatives to have a fourth moment. This will hold when the simulators are used to define sufficiently well-behaved data-generating processes, but will be violated for sufficiently heavy-tailed distributions (e.g. the Cauchy). Finally, (A3) is mild; it holds when the gradient of the log conditional density estimator is Lipschitz continuous, which is for example the case when the conditional density is twice continuously differentiable and the Hessian is bounded (see e.g. Lemma 2.3 of Wright and Recht (2022)). It also holds for models such as mixtures of Gaussians, which are used in mixture density networks and for normalising flows with sufficiently regular transformations; see Table 1 in Liang et al. (2022) for some known Lipschitz continuous transformations. There are two key implications of Theorem˜1. The first is a bound on the variance of the MC objective:
$\text{Var}\left[{\ell}_{\text{MC}}(\phi)\right]\leq\frac{K(\phi)}{n}\left(\left\|G\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}^{4}+1\right),$ (4)
which is obtained by noticing that the bound on $\text{Var}[h_{0}(\phi)]$ is simply a bound on a Monte Carlo objective. From this, we immediately notice that the bound will be large whenever the high-fidelity simulator is expensive and $n$ is small, or whenever the simulator is complex as measured in this Sobolev norm. The second implication is a bound on the variance of the MLMC objective:
$\text{Var}\left[{\ell}_{\text{MLMC}}(\phi)\right]$
$\leq\frac{K_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}^{4}+1\right)+\sum_{l=1}^{L}\frac{K_{l}(\phi)}{n_{l}}\left(\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U}^{m})}+1\right).$ (5)
In order to make this bound small, we need to make each term small. Typically, $\|G^{0}\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}$ will be large, but this will be counter-balanced by taking $n_{0}$ to be large. For the higher-fidelity levels, the number of samples $n_{l}$ will typically be smaller, but this will be counter-balanced by the fact that $\|G^{l}-G^{l-1}\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}$ is small whenever $G^{l-1}$ is a good approximation of $G^{l}$. As we now show, we can directly use ˜5 to get an indication of how to select $n_{0},\ldots,n_{L}$.
@@
Suppose the assumptions of Theorem˜1 hold. Then, the values of $n_{0},\ldots,n_{L}$ which minimise the upper bound on $\text{Var}[{\ell}_{\text{MLMC}}(\phi)]$ in ˜5 given a fixed computational budget; i.e. for $\text{Cost}({\ell}_{\text{MLMC}}(\phi);n_{0},\ldots,n_{L})\leq C_{\text{budget}}$ for some $C_{\text{budget}}>0$ are given by
$n^{\star}_{0}\propto\frac{C_{\text{budget}}}{\sqrt{C_{0}}}\sqrt{\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}^{4}+1},\quad n^{\star}_{l}\propto\frac{C_{\text{budget}}}{\sqrt{C_{l}+C_{l-1}}}\sqrt{\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U}^{m})}+1}.$
See Section˜A.2 for the proof. This result provides useful intuition on how to select the number of sample per level. For instance, consider the case $L=1$, i.e., two levels. If $G^{0}$ is known to be a good approximation of $G^{1}$, it makes sense to allocate a large budget to generating low-fidelity simulations while only a small number of high-fidelity simulations would be sufficient. On the other hand, if $G^{0}$ and $G^{1}$ differ substantially, allocating a larger budget to high-fidelity simulations makes sense, despite the higher cost, as it can help accurately capture this difference. Although Theorem˜2 provides intuition, it may be hard to obtain the optimal $n_{0},\dots,n_{L}$ exactly since it requires computing quantities which are often unknown. For example, computing Sobolev norms can be challenging, and the results implicitly depend on $\phi$ through the constants $K_{0}(\phi),\ldots,K_{L}(\phi)$ of Theorem˜1 (which are unlikely to be tight). This is a common limitation of MLMC theory; see (Giles, 2015, Sec. 2-3).
@@ 5 Numerical Experiments
We compare the performance of our multilevel version of NLE and NPE, termed ML-NLE and ML-NPE, respectively, against their standard counterpart with the MC loss. We use the sbi library (Tejero-Cantero et al., 2020) implementation for NLE and NPE , see Appendix˜B for the details. The code to reproduce our experiments is available at https://github.com/yugahikida/multilevel-sbi.
@@ 5.1 The g-and-k distribution: an illustrative example
We first consider the g-and-k distribution (Prangle, 2020) as an illustrative example. This is a very flexible univariate distribution that is defined via its quantile function and has four parameters, controlling the mean, variance, skewness, and kurtosis respectively, making it challenging for SBI methods. It does not typically have a low-fidelity simulator, so we construct one through a Taylor approximation of the quantile function, see Section˜B.1 for details. This makes it a slightly contrived example, but the fact that the g-and-k allows for an efficient approximation of the likelihood will make it particularly convenient to study the performance of NLE-based methods.
@@
We fix the number of low-fidelity samples ($n_{0}=10^{4}$ for ML-NLE and $n_{0}=10^{3}$ for ML-NPE) and vary the number of high-fidelity samples $n_{1}$ to asses the improvement in performance of our methods as $n_{1}$ increases. For ML-NLE, we compute the Kullback-Leibler divergence (KLD) between the estimated conditional density and a numerical approximation of the likelihood. For ML-NPE, we use the negative log-posterior density (NLPD) of the true $\theta$ under the estimated posterior density as the metric. The results in Figure˜2(a) and 2(b) show that the multilevel versions of NLE and NPE perform better than their standard counterparts (with MC loss) using just a fraction of the high-fidelity samples. Unsurprisingly, the performance of MLE-NLE and ML-NPE improves as $n_{1}$ increases.
In Figure˜2(c), we show an example of the NLE densities learned, with additional examples in Section˜B.1. Our ML-NLE method is able to approximate the almost exact g-and-k density the best. The coverage plot in Figure˜2(d) shows that ML-NPE yields slightly conservative posteriors as opposed to the overconfident posteriors obtained from NPE trained on either all low- or all high-fidelity data.
@@ 5.2 Ornstein–Uhlenbeck process: a popular financial model
Our next experiment involves the Ornstein-Uhlenbeck (OU) process—a stochastic differential equation model commonly used in financial analysis (Minenna, 2003), which in our case outputs a $100$-dimensional Markovian time-series and has three parameters. The process is known to converge to a stationary Gaussian distribution, which we use as the low-fidelity simulator, see Section˜B.3. The example is reproduced from Krouglova et al. (2025), who used it for benchmarking their transfer learning approach to NPE (TL-NPE). TL-NPE first trains an NPE network on low-fidelity simulations, and then uses a small set of high-fidelity data to refine the network parameters until a stopping criterion is met. We implement TL-NPE with $n_{0}=1100$ and $n_{1}=100$, using 10% of data as validation set for their stopping criterion, and compare against ML-NPE with $n_{0}=1000$ and $n_{1}=100$. These values are picked so as to keep the simulation budget same for both the methods.
We compare ML-NPE with TL-NPE for different choices of its hyperparameter in Figure˜3. TL-NPE’s hyperparameter is the number of epochs until the stopping criterion is activated, termed patience, which we let take values in $\{1,20,10^{2},10^{3}\}$. We use two metrics for evaluation: the NLPD of the true $\theta$ under the estimated posterior, and the KLD between the posterior obtained using either TL-NPE or ML-NPE and a reference NPE posterior trained using $n=10,000$ samples. We observe that TL-NPE’s performance is quite sensitive to the choice of the patience value, and the authors of that work do not provide guidance on how to select this in practice (their code fixes this to $20$ with no clear justification). When the number of high-fidelity samples are low ($n_{1}=10$), its performance degrades as patience increases, as a higher patience means TL-NPE is overfitting to the few high-fidelity samples, losing the information gained from training on the low-fidelity data. In contrast, ML-NPE achieves better performance especially when $n_{1}=10$ and does not have any hyperparameter to tune, making it easier to use in practice. We additionally investigate a challenging scenario where the high- and low-fidelity simulators have differing dimensionalities of $\theta$, in which our method exhibits some limitations, possibly due to instability during optimisation; see Section˜B.3.1 for details.
@@
Figure 3: NLPD and KLD for ML-NPE (ours) and TL-NPE (Krouglova et al., 2025) using different patience values (hyperparameter of TL-NPE) over 2020 runs. ML-NPE performs better in the limited high-fidelity data scenario ($n_{1}=10$) (a), whilst being competitive when $n_{1}=100$ (b). TL-NPE shows sensitivity towards its hyperparameter, while ML-NPE does not require any hyperparameter.
Figure 4: MMD ($\downarrow$) across 5000 parameter values for NLE and ML-NLE.
@@ 5.3 Toggle-switch model: a Systems Biology example
We now consider the toggle-switch model (Bonassi et al., 2011; Bonassi and West, 2015). This model describes the interaction between two genes over time, and has seven parameters and a scalar observation at the end of a time interval. Simulators typically use time-discretisation, and the total number of time-steps $T$ acts as a fidelity parameter: running the model with large $T$ incurs larger computational cost but leads to accurate simulations, while smaller $T$ leads to cheap but inaccurate samples. Thus, this model illustrates a setting with more than two fidelity levels (by taking $T$ to be more than two values)—a situation straightforward to handle with our multilevel approach but not with TL-NPE (Krouglova et al., 2025). We take $T_{0}=50$, $T_{1}=80$, and $T_{2}=300$ to be the number of steps for the three fidelity levels with $n_{0}=10^{4}$, $n_{1}=500$, and $n_{2}=100$. Guided by the intuition from Theorem˜2, we allocate a large budget to estimating the difference between the low- and the medium-fidelity simulators, leveraging prior knowledge that this difference is substantial. See Section˜B.4.1 for results under alternative budget allocations. Note that in this case each fidelity level has a different base measure of dimension $2T+1$; however, there is still sufficient seed-matching (see Section˜B.4 for details), and therefore variance reduction, thanks to MLMC. We compare our ML-NLE method with the standard NLE trained on samples from either the low- ($n=12,060$), the medium- ($n=7537$), or the high-fidelity simulator ($n=2010$). The number of training data $n$ in each case is selected so as to match the total computational cost of simulation between ML-NLE and NLE, see Section˜B.4. Figure˜4 reports the maximum mean discrepancy (MMD) (Gretton et al., 2012) between $500$ samples from the learned conditional densities and $500$ samples from the high-fidelity simulator across for $5000$ different parameter values. We observe that ML-NLE performs better than all the NLE baselines for the same computational cost.